### v1

# Evaluating Memory in LLM Agents via Incremental Multi-Turn Interactions

Abstract
Recent benchmarks for Large Language Model (LLM) agents primarily focus on evaluating reasoning, planning, and execution capabilities, while another critical component—memory, encompassing how agents memorize, update, and retrieve long-term information—is under-evaluated due to the lack of benchmarks. We term agents with memory mechanisms as memory agents. In this paper, we identify four core competencies essential for memory agents: accurate retrieval, test-time learning, long-range understanding, and conflict resolution. Existing datasets either rely on limited context lengths or are tailored for static, long-context settings like book-based QA, which do not reflect the interactive, multi-turn nature of memory agents that incrementally accumulate information. Furthermore, no existing benchmarks cover all four competencies. Therefore, we introduce MemoryAgentBench, a new benchmark specifically designed for memory agents. Our benchmark combines reformulated existing datasets with newly constructed ones, covering the above four memory competencies, providing a systematic and challenging testbed for assessing memory quality. We evaluate a diverse set of memory agents, ranging from simple context-based and retrieval-augmented generation (RAG) systems to advanced agents with external memory modules and tool integration. Empirical results reveal that current methods fall short of mastering all four competencies, underscoring the need for further research into comprehensive memory mechanisms for LLM agents.
1 Introduction
Large Language Model (LLM) agents have rapidly transitioned from proof-of-concept chatbots to end-to-end systems that can write software [42], control browsers [34], and reason over multi-modal inputs. Frameworks such as Manus, OWL [16], OpenHands [42], and Codex routinely solve complex, tool-rich tasks and achieve state-of-the-art results on agentic benchmarks like GAIA [31] and SWE-Bench [18]. Yet these evaluations focus almost exclusively on reasoning (planning, tool using, code synthesis) and leave the equally important question of memorization (abstraction, storing, updating, retrieving) largely under-explored.
Recent memory-centric architectures—ranging from parametric memory systems like MemoryLLM [44], SELF-PARAM [43], and M+[46] to commercial token-level memory solutions such as MemGPT[36, 27], Mem0[7], Cognee, and Zep[37]—employ diverse strategies for storing and retrieving past information. Despite growing interest, their real-world effectiveness remains largely anecdotal, and there is currently no unified benchmark for systematically evaluating the quality of memory in agents. In this paper, we refer to agents equipped with memory mechanisms as Memory Agents, where memory can take various forms, including parameters, vectors, textual histories, or external databases. In this paper, we primarily focus on memory agents that utilize textual histories and external databases, as these approaches are most commonly deployed in real-world applications. In contrast, memory encoded in model parameters [44, 46, 51] remains largely within academic research and is typically less capable than proprietary memory systems equipped on closed-sourced API models.
To evaluate memory agents, we identify four complementary competencies (Examples shown in Figure 1): (1) Accurate Retrieval: The ability to extract the correct snippet in response to a query. This can involve one-hop or multi-hop retrieval, as long as the relevant information can be accessed with a single query. (2) Test-Time Learning: The capacity to incorporate new behaviors or acquire new skills during deployment, without additional training. (3) Long-Range Understanding: The ability to integrate information distributed across extended contexts ( 100k tokens) and answer questions requiring a global understanding of the entire sequence. (4) Conflict Resolution: The skill to revise, overwrite, or remove previously stored information when faced with contradictory evidence, aligning with goals in model editing and knowledge unlearning tasks [30, 45].
Previous datasets developed to evaluate memory in language models have notable limitations. Early benchmarks such as LOCOMO [29] ( 9k tokens), LooGLE[23] ( 24k tokens), and LongBench[3] ( 20k tokens) feature relatively short contexts that no longer challenge current models. More recent datasets like NovelQA[40] (200k tokens), NOCHA[19] (127k tokens), Loong[41] (100k tokens), and -Bench[53] (150k tokens) extend the context length to evaluate global reasoning and retrieval capabilities. However, these datasets were primarily designed for evaluating long-context language models rather than memory agents. The reason that long-context benchmarks cannot be directly used to evaluate memory agents is as follows. There is a fundamental distinction between memory and long context: memory serves as a compressed and distilled representation of past information. Rather than storing all historical content verbatim, memory selectively extracts salient details, removes irrelevant information, and often incorporates new inferences derived from prior experiences. Consequently, memory agents are designed to process context incrementally—absorbing input piece by piece, abstracting and consolidating information over time, generating new inferences, and learning novel rules from accumulated history. For this reason, datasets that provide the entire context in a single block are not directly applicable to evaluating memory agents. A more recent effort, LongMemEval [48], seeks to address this limitation by using synthetic long-form conversations, which can be injected into memory gradually, session by session. Nonetheless, its evaluation framework remains constrained by limited topical diversity and less realistic interaction patterns, reducing its applicability to real-world memory agent scenarios.
To address these limitations, we introduce a unified benchmark framework, MemoryAgentBench, specifically designed to evaluate a broad spectrum of memory mechanisms in agent systems. We also provide a framework for memory agent evaluation. In this framework, agents are presented with sequences of textual inputs that simulate multi-turn interactions with users. We repurpose existing datasets originally developed for long-context LLM evaluation by segmenting their inputs into multiple chunks and feeding them incrementally to the agent. However, since these datasets do not fully capture all four targeted memory competencies, we also introduce two new datasets: EventQA and FactConsolidation, designed to evaluate accurate retrieval and conflict resolution, respectively. Our benchmark includes evaluations of state-of-the-art commercial memory agents (such as Mem0 and MemGPT), long-context agents that treat the full input as memory, and RAG agents that extend their memory through retrieval methods. We examine how techniques developed for long-context models and RAG transfer to the memory agent setting, and how commercial memory agents perform under more challenging, competency-specific tests. By providing a consistent evaluation protocol across diverse agent architectures and datasets, MemoryAgentBench delivers comprehensive insights into agent performance across the four core memory competencies.
Our contributions are summarized as follows:
Datasets: We re-structure existing datasets and create two new datasets to construct a comprehensive benchmark, covering four distinct memory competencies.
Framework: We provide a unified evaluation framework, and open-source the codebase and datasets to encourage reproducibility and further research.
Empirical Study: We implement various simple agents with diverse memory mechanisms, adopt commercial agents, and evaluate these agents on our proposed benchmark. With our results, we show that existing memory agents, while effective in some tasks, still face significant challenges on some aspects.
2 Related Work
2.1 Benchmarks with Long Input
In this section, we review prior work on long-context benchmarks. Early benchmarks designed for long-context evaluation include LongBench[3] and LooGLE[23], with average input lengths of approximately 20k and 24k tokens, respectively. More recent benchmarks—such as -Bench [53], HELMET[50], RULER[15], NOCHA[19], NoLiMa [33] and LongBench V2[4]—extend context lengths to over 100k tokens and are primarily intended to evaluate the capabilities of long-context models. However, despite their scale, these benchmarks are not designed to assess memory agents, and no prior work has repurposed them for that goal. More recently, LOCOMO [29] and LongMemEval [48] have been proposed specifically for evaluating memory agents. While promising, LOCOMO still features relatively short conversations (9k), and LongMemEval uses synthetic conversations with limited topical diversity, making the dialogues less realistic and potentially less representative of real-world memory use cases.
2.2 Agents with Memory Mechanisms
Memory mechanisms are attracting more and more attention lately [47]. Recent advancements in LLMs have demonstrated the capability to process extended context lengths, ranging from 100K to over 1 million tokens. For instance, models such as GPT-4o [35] and Claude 3.7 [1] can handle inputs of approximately 100K to 200K tokens, while models like Gemini 2.0 Pro [8] and the GPT-4.1 series extend this capacity beyond 1 million tokens. These strong long-context capabilities enable a simple yet effective form of memory: storing information directly within the context window. However, this approach is inherently constrained by a hard limit—once the context window is exceeded, earlier information must be discarded.
In parallel, RAG continues to serve as a dominant paradigm for managing excessive context. By retrieving relevant information from earlier context and feeding it to the LLM, RAG allows systems to overcome context length limitations. For example, OpenAI’s recent memory functionality111https://openai.com/index/memory-and-new-controls-for-chatgpt/ combines explicit user preference tracking with retrieval-based methods that reference prior interactions. RAG methods can be broadly classified into three categories: 1. Simple RAG: These methods rely on string-matching techniques such as TF-IDF, BM25 [38], and BMX [25], which are entirely non-neural and operate on string-level similarity. 2. Embedding-based RAG: This class leverages neural encoders, primarily transformers, to map text into dense vector representations [49]. Early methods like DPR [20] and Contriever [17] are based on BERT [9], while more recent models such as NV-Embed-v2 [22] utilize decoder-only backbones and achieve significantly improved retrieval performance. 3. Structure-Augmented RAG: These approaches enhance retrieval with structural representations such as graphs or trees. Representative systems include GraphRAG [10], RAPTOR [39], HippoRAG-V2 [12], Cognee, Zep [37], and Mem0 [7], where Mem0 also offers a graph-augmented variant, Mem0g, built on structured factual knowledge. Despite their effectiveness, RAG-based methods face challenges with ambiguous queries, multi-hop reasoning, and long-range comprehension. When questions require integrating knowledge across an entire session or learning from long, skill-encoding inputs, the retrieval mechanism—limited to the top-k most relevant passages—may fail to surface the necessary information. To address these limitations, Agentic Memory Agents introduce an iterative, decision-driven framework. Rather than relying on a single-pass retrieval, these agents dynamically process the query, retrieve evidence, reflect, and iterate through multiple retrieval and reasoning cycles. Examples include MemGPT [36], Self-RAG [2], and Auto-RAG [52]. This agentic design is particularly effective for resolving ambiguous or multi-step queries. Nonetheless, these methods remain fundamentally constrained by the limitations of RAG—namely, the inability to fully understand or learn from long-range context that is inaccessible via retrieval alone.
3 MemoryAgentBench
3.1 Aspects of the Evaluation
The evaluation of memory agents encompasses the following key dimensions:
Accurate Retrieval (AR)
The task of accurately retrieving information has been extensively explored in prior work. In the domain of long-context modeling, the Needle-in-a-Haystack (NIAH) task is widely used to evaluate a model’s ability to locate a specific value based on a given key within a lengthy input. Extensions such as multi-value NIAH further test the model’s capacity to retrieve multiple values scattered across the input context. In the RAG setting, this corresponds to document-based QA, where the model must identify and extract relevant snippets from one or more documents to answer a query. These snippets might reside in a single location or be distributed across multiple documents. In this paper, we focus on agentic settings, where the “long context” or “multiple documents” become long-form conversations. We define Accurate Retrieval (AR) as the ability of an agent to identify and retrieve important information that may be dispersed throughout a long dialogue history.
Test-Time Learning (TTL)
An essential capability for real-world agents is the ability to acquire new skills dynamically through interaction with users. This mirrors the concept of In-Context Learning (ICL) in LLMs, where the model learns from a prompt containing a small number of examples, often framed as few-shot classification tasks. Ideally, performance improves with additional examples in the prompt. In the conversational agent setting, prompts are replaced by dialogue histories. We define Test-Time Learning (TTL) as the agent’s ability to learn to perform new tasks directly from the conversation. This property is crucial for enabling self-evolving agents that can continuously adapt and improve in real-world deployments.
Long-Range Understanding (LRU)
Long-range understanding refers to the agent’s ability to form abstract, high-level comprehension over extended conversations. For example, when a user narrates a long story, the agent should retain the content and derive a holistic understanding rather than just recall isolated facts. We define Long-Range Understanding (LRU) as the ability to reason about long-form inputs and answer high-level questions that require an understanding of the overall content, rather than detailed recall. An example question might be: “Summarize the main experiences of Harry Potter.”
Conflict Resolution (CR)
In long-term interactions, agents often face evolving or conflicting information—whether about the external world (e.g., changes in political leadership) or user-specific facts (e.g., a new occupation). This challenge is closely related to model editing [30, 11] and knowledge unlearning [45], which focus on modifying or removing factual knowledge from language models. We define Conflict Resolution (CR) as the agent’s ability to detect and resolve contradictions between existing knowledge and newly acquired information, ensuring the agent remains aligned with current realities and user states. CR is distinct from Abstractive Retrieval (AR) in two key ways. (1) Certain questions requiring CR cannot be answered solely through AR. As illustrated in Figure 1, an agent that retrieves all facts related to pears may fail to identify the updated information in the second message. (2) In AR, earlier messages remain relevant and should be retained, even when multiple pieces of evidence are required. In contrast, CR involves identifying outdated or incorrect information and discarding it. That is, AR requires preservation of all related content, whereas CR requires overwriting prior facts to reflect the most up-to-date truth.
| Capability | Benchmarks / Tasks | # of Sequences | # of QAs | Avg Len |
| Accurate Retrieval | RULER-QA [15] | 2 | 200 | 309K |
| RULER-NIAH-MQ [15] | 1 | 100 | 448K | |
| Bench-QA [53] | 100 | 100 | 183K | |
| LongMemEval (S*) [48] | 5 | 300 | 355K | |
| EventQA (ours) | 5 | 500 | 534K | |
| Test-Time Learning | BANKING-77 | 1 | 100 | |
| CLINC-150 | 1 | 100 | ||
| NLU | 1 | 100 | 103K | |
| TREC (Coarse) | 1 | 100 | ||
| TREC (Fine) | 1 | 100 | ||
| Movie-Rec Redial [14] | 1 | 200 | 1.44M | |
| Long-Range Understanding | Bench-Sum [53] | 100 | 100 | 172K |
| Conflict Resolution | FactConsolidation-SH (ours) | 1 | 100 | 262K |
| FactConsolidation-MH (ours) | 1 | 100 |
3.2 Dataset Preperation
In this section, we describe how we adopt existing datasets and construct new ones for evaluating each aspect introduced in Section 3.1. All datasets with their categories are shown in Table 1. We introduce the details in datasets curation in Appendix A.
Datasets for Accurate Retrieval (AR)
We adopt five datasets to evaluate the accurate retrieval capability of memory agents. Four are adapted from existing benchmarks, and one is newly constructed: (1) RULER-QA: This is a NIAH-style QA task where a long passage contains single (QA-1) or multiple (QA-2) snippets answering the input question. The agent must identify and extract relevant snippets from the extended context. (2) NIAH-MQ: We use the multiple-query (MQ) version of the NIAH dataset from RULER [15], where each query seeks a different numeric value embedded in a long passage. The agent must accurately retrieve multiple distinct answers. (3) Bench-En.QA: This task from Bench presents free-form QA questions based on entire books, with all entities replaced by fictitious names to avoid contamination from model pretraining. Compared to synthetic datasets like RULER-QA, this benchmark is more realistic and challenging due to the natural narrative structure of books. (4) LongMemEval: This benchmark evaluates memory agents on long dialogue histories. Although task types like information extraction (IE) or multi-session reasoning are included, most tasks can be reformulated as single-retrieval problems requiring agents to retrieve the correct segments spanning a long multi-turn conversation. Among these, LongMemEval is already formatted for agent-based evaluation with session separation. We use the original LongMemEval(S) dataset (110K tokens) and reformulated chat history into five long dialogues (355K tokens) with 300 questions (LongMemEval (S*) in Table 1). We create LongMemEval (S*) specifically for increasing the number of questions per context, mitigating the exhaustive needs of reconstructing the memory for each quesiton. (5) EventQA (ours): We introduce EventQA this reasoning style NIAH task to evaluate agents’ ability to recall and reason about temporal sequences in long-form narratives. In this dataset, the agent is required to read a novel and select the correct event from a series of candidates after receiving up-to five previous events. For these datasets, which are originally designed for long-context modeling, we split documents into chunks and sequentially inject them into the agent.
Datasets for Test-Time Learning (TTL)
We evaluate TTL via two task categories: (1) Multi-Class Classification (MCC): We adopt five classification datasets used in prior TTL work [5, 50]: BANKING77 [6], CLINC150 [21], TREC-Coarse, TREC-Fine [26], and NLU [28]. Each task requires the agent to map sentences to class labels, leveraging previously seen labeled examples in context. (2) Recommendation (Recom): We use the Redial [24] dataset to evaluate movie recommendation via dialogue. Following the setup from He et al. [13], the agent is exposed to thousands of movie-related dialogue turns and is asked to recommend twenty relevant movies based on the long interaction history.
Datasets for Long Range Understanding (LRU)
We adopt the Summarization task En.Sum from -Bench [53]. The agent is required to analyze and organize the plot and characters of the novel, and then compose a summary of 1000 to 1200 words.
Datasets for Conflict Resolution (CR)
To assess whether an agent can consolidate conflicting factual updates and reason over them, we construct a new dataset called FactConsolidation. Specifically, We build this benchmark using counterfactual edit pairs from MQUAKE [54]. Each pair contains a true fact and a rewritten, contradictory version. These are ordered such that the rewritten (new) fact appears after the original, simulating a realistic update scenario. We concatenate multiple such edit pairs to create long contexts of length 32K, 64K, 262K. We then adpot MQUAKE’s original questions and categorize them into: (1) FactConsolidation-SH (Ours) (SH means Single-Hop), requiring direct factual recall (e.g., “Which country was tool created in?”), and (2) FactConsolidation-MH (Ours) (MH refers to Multi-Hop), requiring inference over multiple facts (e.g., “What is the location of death of the spouse of person ?”). Agents are prompted to prioritize later information in case of conflict and reason based on the final memory state. This setup directly evaluates the strength and consistency of conflict resolution over long sequences.
3.3 Different Categories of Memory Agents
We evaluate three major types of memory agents that reflect common strategies for handling long-term information: Long-Context Agents, RAG Agents, and Agentic Memory Agents. These approaches differ in how they store, retrieve, and reason over past inputs.
Long Context Agents
Modern language models often support extended context windows ranging from 128K to over 1M tokens. A straightforward strategy for memory is to maintain a context buffer of the most recent tokens. For example, in a model with a 128K-token limit, the agent concatenates all incoming chunks until the total exceeds the window size. Once the limit is reached, the earliest chunks are evicted in a FIFO (first-in, first-out) manner. This agent design relies solely on positional recency and assumes the model can attend effectively over the current context window.
RAG Agents
RAG-based agents address context limitations by storing past information in an external memory pool and retrieving relevant content as needed. We consider three RAG variants: (1) Simple RAG Agents: All input chunks are stored as raw text. During inference, a keyword or rule-based string matching mechanism retrieves relevant passages. (2) Embedding-based RAG Agents: Each input chunk is embedded and saved. At query time, the agent embeds the query and performs retrieval using cosine similarity between embeddings. (3) Structure-Augmented RAG Agents: After ingesting all input chunks, the agent constructs a structured representation (e.g., knowledge graph or event timeline). Subsequent queries are answered based on this structured memory.
Agentic Memory Agents
Agentic memory agents extend beyond static memory stores by employing agentic loops—iterative reasoning cycles in which the agent may reformulate questions, perform memory lookups, and update its working memory. These agents are designed to simulate a more human-like process of recalling, verifying, and integrating knowledge.
3.4 Datasets and Agents Formulation
Datasets Formulation
We standardize all datasets into the format: (chunks), (questions), and (answers), where denotes the -th chunk wrapped to construct a user message with instructions of memorizing the content in a sequential input, and represents a single conversation. Each chunk is accompanied by instructions prompting the agent to memorize its contents. Example prompts are provided in Appendix B.1.
When curating datasets like EventQA and FactConsolidation, we deliberately design scenarios where multiple questions follow a single context. This allows us to probe the model’s memory multiple times with one sequential injection. For example, in LME (S*), five contexts are paired with 300 questions (shown in Table 1). This design choice reflects a key trend: as LLMs support increasingly long context windows and memory agents become more capable of handling extended inputs, evaluation datasets must also scale accordingly. Injecting 1M tokens for just one question is resource-inefficient, whereas associating the same input with many questions provides significantly higher utility.
Agents Formulation
In our framework, all agents are required to take the chunks one by one, absorb them into memory, and incrementally update the memory. After seeing all the chunks, we ask the agent to answer the related questions.
4 Experiments
4.1 Experimental Setup
The datasets are split into four categories and the statistics of all datasets are also shown in Table 1. The evaluation metrics for all datasets are shown in Table 5 in Appendix A, along with more dataset details. Then for the agents, as described in Section 3.3, we consider three categories of agents: Long-Context Agents, RAG agents and Agentic Memory Agents, where RAG Agents can be further split into Simple RAG Agents, Embedding-based RAG Agents and Structure-Augmented RAG Agents.
For chunk size settings, we choose a chunk size of 512 for the RULER-QA, NIAH-MQ, and LME(S*) tasks in AR, as well as for all tasks in CR. This is mainly because these tasks are composed of long texts synthesized from multiple short texts. For other tasks, we use a chunk size of 4096. Considering computational overhead, we uniformly use a chunk size of 4096 for the Mem0 and Cognee methods. We report the settings of the chunk size in Table 15 in Appendix C.
| AR | TTL | LRU | CR | |||||||
| Agent Type | RULER-QA | NIAH-MQ | Bench-QA | LME(S*) | EventQA | MCC | Recom | Bench-Sum | FactCon-SH | FactCon-MH |
| Long-Context Agents | ||||||||||
| GPT-4o | 61.5 | 25.0 | 55.4 | 32.0 | 77.2 | 87.6 | 12.3 | 32.2 | 60.0 | 5.0 |
| GPT-4o-mini | 53.5 | 22.8 | 44.9 | 30.7 | 59.0 | 82.4 | 15.1 | 28.9 | 45.0 | 5.0 |
| GPT-4.1-mini | 74.5 | 94.8 | 45.8 | 55.7 | 82.6 | 75.6 | 16.7 | 41.9 | 36.0 | 5.0 |
| Gemini-2.0-Flash | 73.0 | 83.8 | 53.2 | 47.0 | 67.2 | 84.0 | 8.7 | 23.9 | 30.0 | 3.0 |
| Claude-3.7-Sonnet | 65.0 | 38.0 | 50.6 | 34.0 | 74.6 | 89.4 | 18.3 | 52.5 | 43.0 | 2.0 |
| GPT-4o-mini | 53.5 | 22.8 | 44.9 | 30.7 | 59.0 | 82.0 | 15.1 | 28.9 | 45.0 | 5.0 |
| Simple RAG Agents | ||||||||||
| BM25 | 61.0 | 100.0 | 45.6 | 45.3 | 74.6 | 75.4 | 13.6 | 20.9 | 56.0 | 3.0 |
| Embedding RAG Agents | ||||||||||
| Contriever | 26.5 | 2.5 | 38.1 | 15.7 | 66.8 | 70.6 | 15.2 | 21.2 | 18.0 | 7.0 |
| Text-Embed-3-Small | 52.0 | 7.2 | 44.4 | 48.3 | 63.0 | 70.0 | 15.3 | 25.7 | 28.0 | 3.0 |
| Text-Embed-3-Large | 49.0 | 19.5 | 50.1 | 52.3 | 70.0 | 72.4 | 16.2 | 21.6 | 28.0 | 4.0 |
| NV-Embed-v2 | 83.0 | 73.5 | 51.4 | 55.0 | 72.8 | 69.4 | 13.5 | 20.7 | 55.0 | 6.0 |
| Structure-Augmented RAG Agents | ||||||||||
| RAPTOR | 33.5 | 15.8 | 31.3 | 34.3 | 45.8 | 59.4 | 12.3 | 13.4 | 14.0 | 1.0 |
| GraphRAG | 47.0 | 38.3 | 35.8 | 35.0 | 34.4 | 39.8 | 9.8 | 0.4 | 14.0 | 2.0 |
| HippoRAG-v2 | 71.0 | 67.5 | 45.7 | 50.7 | 67.6 | 61.4 | 10.2 | 14.6 | 54.0 | 5.0 |
| Mem0 | 28.0 | 4.8 | 22.4 | 36.0 | 37.5 | 3.4 | 10.0 | 0.8 | 18.0 | 2.0 |
| Cognee | 33.5 | 4.0 | 19.7 | 29.3 | 26.8 | 35.4 | 10.1 | 2.3 | 28.0 | 3.0 |
| Agentic Memory Agents | ||||||||||
| Self-RAG | 38.5 | 8.0 | 28.5 | 25.7 | 31.8 | 11.6 | 12.8 | 0.9 | 19.0 | 3.0 |
| MemGPT | 39.5 | 8.8 | 20.8 | 32.0 | 26.2 | 67.6 | 14.0 | 2.5 | 28.0 | 3.0 |
4.2 Overall Performance Comparison
Table 2 presents the overall performance across different benchmarks. We summarize the key findings as follows: (1) Superiority of RAG methods in Accurate Retrieval Tasks. Most RAG Agents are better than the backbone model “GPT-4o-mini” in the tasks within the Accurate Retrieval Category. This matches our intuition where RAG agents typically excel at extracting a small snippet of text that is crucial for answering the question. (2) Superiority of Long-Context Models in Test-Time Learning and Long-Range Understanding. Long-context models achieve the best performance on TTL and LRU. This highlights a fundamental limitation of RAG methods and commercial memory agents, which still follow an agentic RAG paradigm. These systems retrieve only partial information from the past context, lacking the ability to capture a holistic understanding of the input—let alone perform learning across it. (3) Limitation of All Existing Methods on Conflict Resolution. Although being a well-discussed task in model-editing community [32, 11], resolving conflict poses a significant challenge on memory agents. We observe that all methods fail on the multi-hop situation (with achieving at most 6% accuracy). Only long context agents can achieve fairly reasonable results on single-hop scenarios. In Section 4.4.2, we show that current reasoning models can have much better performance, while it does not change the conclusion that Conflict Resolution still poses a significant challenge on all memory mechanisms. (4) Limited Performance of Commercial Memory Agents. Commercial memory agents such as MemGPT and Mem0 exhibit limited performance across a broad range of benchmarks. This shortfall can be attributed to three primary factors. First, these systems frequently fail to capture and preserve sufficient information when storing inputs into memory. For example, Mem0 depends on extracting factual knowledge from inputs, an approach that inherently discards a substantial portion of the original content. As a result, reconstructing inputs and supporting downstream tasks such as question answering becomes significantly more challenging. While Mem0 has demonstrated relatively strong performance on conversational tasks such as LOCOMO—where information density is comparatively low—it tends to perform poorly on benchmarks containing dense informational content, including RULER and -Bench. For tasks emphasizing Time-to-Live (TTL) and Least Recently Used (LRU) retrieval, these limitations are often even more pronounced. Second, both MemGPT and Mem0 rely on retrieval mechanisms that only access a subset of stored information. In the case of Mem0, retrieval is typically performed a single time, similar to conventional RAG methods, constraining the breadth of information available for reasoning. MemGPT, although adopting a more agentic framework that permits multiple retrieval iterations, does not maintain temporal or structural metadata about the stored content. Consequently, the agent is unable to reconstruct longer documents in their original form, which adversely affects performance on LRU and other tasks requiring structured memory retrieval. Finally, methods such as MemGPT depend heavily on embedding-based retrieval mechanisms, which can be insufficient for fine-grained tasks like NIAH, where locating specific, precise information (“the needle in the haystack”) is essential. These embedding-based approaches often struggle to distinguish subtle contextual nuances that are critical for accurate retrieval in such settings.
4.3 Ablation Study
In this Section, we conduct experiments and result analysis along four dimensions: Input Chunk Size, Retrieval TopK, Validation of Dataset, and Computational Latency. More detailed experimental results are provided in Appendix C.
4.4 Ablation Study on Input Chunk Size
To understand how chunk size impacts performance, particularly for RAG methods and agentic memory agents, we conduct an additional analysis where we vary the chunk size while fixing the number of retrieved chunks to 10. The results are presented in Figure 2. From the figure, we observe the following: (1) In the RULER-QA task, reducing chunk size has little effect on BM25 performance. This is expected, as BM25 relies on term-frequency-based scoring and document-level ranking, and does not inherently benefit from finer-grained segmentation beyond the impact on term distributions. In contrast, embedding-based methods—including MemGPT, which uses text-embedding-3-small as its retriever—consistently perform better with smaller chunks. This suggests that finer segmentation improves the granularity and relevance of retrieved results for models that rely on dense semantic representations. (2) In Bench-Sum, however, smaller chunk sizes lead to worse performance. This task requires the agent to summarize an entire conversation, and smaller chunks correspond to fewer available tokens per retrieval. As a result, the agent has access to less context, which degrades summarization quality. The results suggest that, when resources permit, using smaller chunk sizes and increasing the number of retrieval calls during memory construction can improve performance on Accurate Retrieval (AR) tasks. Finer-grained segmentation enhances the relevance of retrieved information, particularly for embedding-based methods. However, for tasks requiring Long-Range Understanding (LRU), varying the chunk size hurts the performance. This is likely because RAG methods are inherently less suited for tasks that demand integration of information across a large, coherent context.
4.4.1 Ablation Study on Retrieval TopK
In our experiments, although we report results with the number of retrieved chunks set to 10 in Table 2, we also conducted ablation studies with varying retrieval sizes. A subset of these results is visualized in Figure 3, with the full results provided in Table 11. The results indicate that increasing the number of retrieved chunks generally improves performance across most tasks. It is worth noting that, with a chunk size of 4096 tokens, retrieving 10 chunks already yields an input of approximately 40k tokens. This places significant demands on model capacity. Due to this high token volume, we do not evaluate settings with 20 retrieved chunks.
| FactCon-SH | FactCon-MH | |||
| 6K | 32K | 6K | 32K | |
| GPT-4o | 92.0 | 88.0 | 28.0 | 10.0 |
| O4-mini | 100.0 | 61.0 | 80.0 | 14.0 |
4.4.2 Validation of Dataset FactConsolidation
As the performance of different models on this dataset remains drastically low, we turn to the stronger reasoning model o4-mini and validate our dataset by checking the performance of o4-mini on a smaller version of this dataset. The results are shown in Table 3.
| 512 | 4096 | |||
| M.C. | Q.E. | M.C. | Q.E. | |
| GPT-4o-mini | 0.09 | 5.2 | 0.07 | 5.1 |
| BM25 | 0.11 | 0.79 | 0.10 | 1.8 |
| Contriever | 7.2 | 0.76 | 1.7 | 2.0 |
| Text-Embed-3-Large | 6.3 | 0.54 | 5.4 | 1.8 |
| NV-Embed-v2 | 93.4 | 0.83 | 42.9 | 1.8 |
| RAPTOR | 151 | 0.51 | 133 | 0.60 |
| GraphRAG | 123 | 9.9 | 90.4 | 9.4 |
| HippoRAG-v2 | 817 | 1.1 | 284 | 2.6 |
| Mem0 | 14644 | 1.2 | 2140 | 1.2 |
| Cognee | 8309 | 33.2 | 962 | 4.5 |
| Self-RAG | 8.4 | 2.0 | 6.7 | 1.7 |
| MemGPT | 413 | 10.6 | 93.3 | 11.4 |
4.4.3 Analysis of Computational Latency
To illustrate the latency of various memory agents in terms of (1) Memory Construction (M.C.); (2) Query Execution (Q.E.), we randomly choose 5 examples from RULER-QA2 and LME (S*), and report the latency of various memory agents. This part of experiments is done on a server with Four NVDIA L40 GPU and AMD EPYC 7713 64-Core CPU. We use the NV-Embed-v2 (7B) as the embedding model in HippoRAG-v2. We show the summarized results in Table 4 and put the full results in Table 12 and 14. From the table, we find that using a smaller chunk size requires significantly more time for memory construction, especially for methods such as HippoRAG-v2, Mem0, Cognee, and MemGPT. Meanwhile, methods such as Mem0, Cognee need extremely high resources when constructing the memory, which may pose challenges in real-world applications.
5 Conclusion and Future Work
In this paper, we introduce MemoryAgentBench, a unified benchmark designed to evaluate memory agents across four essential competencies: accurate retrieval, test-time learning, long-range understanding, and conflict resolution. While prior benchmarks focus largely on skill execution or long-context reasoning, MemoryAgentBench fills a critical gap by assessing how agents store, update, and utilize long-term information across multi-turn interactions. To build this benchmark, we restructure existing datasets and propose two new ones—EventQA and FactConsolidation—tailored to stress specific memory behaviors often overlooked in prior work. We evaluate a wide spectrum of agents, including long-context models, RAG-based systems, and commercial memory agents, under a consistent evaluation protocol. Our results reveal that, despite recent advances, current memory agents still exhibit substantial limitations when faced with tasks requiring dynamic memory updates and long-range consistency. One limitation of our work is that the datasets used in MemoryAgentBench are primarily synthetic, which may not fully reflect the characteristics of real-world user conversations. As future work, we aim to collect and curate more realistic, real-world datasets aligned with our four competencies to further enrich and diversify the benchmark and provide more comprehensive evaluations for memory agents.
Acknowledgment
We thank Kevin Lin for engaging in thoughtful discussions around the overall idea and latency evaluation. His input played a role in shaping the evaluation pipeline for the memory agents.
Appendix A Details of Dataset
Here we provide a detailed introduction to the datasets used for evaluating the four core competencies, including the dataset curation, corresponding metrics, average context length, and a brief description. Details are shown in Table 5.
A.1 Accurate Retrieval (AR)
We use five datasets to evaluate the accurate retrieval capability of memory agents.
(1) RULER-QA
We adopt two QA datasets from [15]. These datasets provide multiple synthetic contexts of varying lengths, ranging from 3K to over 200K tokens. We select 100 questions from the datasets with shorter contexts. For each of these 100 questions, we collect all the documents of them, removed duplicate content, and then shuffled and concatenated them to create new long contexts of 197K or 421K tokens, making sure the new contexts containing the gold passage. Since most answers are short informational entities, such as years, names, or yes/no responses, we use substring exact match (SubEM) as the evaluation metric. SubEM measures whether the predicted answer exactly matches the gold answer as a substring, which is a common standard in question answering systems.
(2) NIAH-MQ
We choose a context with a length of 448K tokens, containing 100 groups and a total of 400 queries. We first check whether these queries appeared in the context. Then, we randomly shuffl the queries and their corresponding numbers evenly to prevent them from clustering together in the long context. The primary evaluation criterion is whether the agent can successfully retrieve the correct numbers. Therefore, we use average recall as the evaluation metric.
(3) Bench-En.QA
We borrow this dataset from [53]. It is a QA task using novels in which character names have been replaced. This makes the content more coherent and closer to real-world scenarios. We use ROUGE F1 for this dataset since answers are mostly entity names.
(4) LongMemEval
This is a dialogue-based QA dataset. For LME(S*), we use multiple historical conversation data segments, arrange them in chronological order, and finally concatenate them to create five long conversation histories, each with a length of approximately 355K tokens. Since some of the questions have open-ended answers, we adopt the approach used in previous work and employ the GPT-4o model to assess whether the agent’s responses meet the requirements. If a response is deemed satisfactory, it is marked as True. Finally, we calculate the proportion of satisfactory responses as the evaluation metric.
(5) EventQA
Using five books from Bench (each >390K tokens, counted using the gpt-4o-mini tokenizer), we identify the ten most frequently mentioned characters via SpaCy NER. We extract 101 events experienced by key characters using gpt-4o. For each event, we construct a 6-way multiple-choice question by pairing the true event with five distractors generated via gpt-4o. The agent receives up-to five previous events and must identify the correct continuation. We report the mean accuracy over 100 such questions per book, and ultimately present the average accuracy across all five books.
| Category | Dataset | Metrics | Avg. Length | Description |
| Accurate Retrieval | RULER-QA1 | SubEM | 197K | Gold passage retrieval QA. |
| RULER-QA2 | 421K | |||
| RULER-NIAH-MQ | Recall | 448K | Retrieve multiple “needles” from the “haystack”. | |
| Bench-QA | ROUGE F1 | 183K | Novel QA with entity replacement. | |
| LongMemEval (S) | Model Based Acc. | 110K | Dialogues based QA. | |
| LongMemEval (S*) | 355K | |||
| EventQA (ours) | Accuracy | 534K | Novel multiple-choice QA on characters events. | |
| Test-time Learning | BANKING77 | Accuracy | 103K | Banking intent classification, 77 labels |
| CLINC150 | Intent classification, 151 labels | |||
| NLU | Task intent classification, 68 labels | |||
| TREC Coarse | Question type classification, 6 labels | |||
| TREC Fine | Question type classification, 50 labels | |||
| Movie Recommendation | Recall@5 | 1.44M | Recommend movies based on provided dialogues examples. | |
| Long Range Understanding | ||||
| Bench-Sum | Model Based F1 | 172K | Novel summarization with entity replacement. | |
| Conflict Resolving | FactConsolidation-SH (ours) | SubEM | 262K | Conflict solving in single hop reasoning. |
| FactConsolidation-MH (ours) | Conflict solving in multiple hop reasoning. |
A.2 Test-time Learning (TTL)
[…]
A.3 Long-Range Understanding (LRU)
We evaluate LRU via the Summarization task En.Sum from -Bench [53]. We follow the settings from [50] and use the GPT-4o model in evaluating the summarized text. In this process, we assess the fluency of the input text (scored as 0 or 1) and use the dot product of this score with the F1 score as the final evaluation metric.
A.4 Conflict Resolution (CR)
We use counterfactual edit pairs from the MQUAKE [54] dataset. Each sentence containing information was assigned a number. For each edit pair, the sentence representing outdated information (the distractor) is given a smaller number, while the sentence representing more recent information (the one containing the answer) is given a larger number. We then concatenate these sentences into a long context in order according to their assigned numbers. We evaluate the CR via two datasets: Single-Hop Editing and Multi-Hop Editing. In these tasks, the agent’s responses are mostly informational entities. Therefore, we also use SubEM (Substring Exact Match) as the evaluation metric.
Appendix B Prompts
We introduce some example prompts used in this section.
B.1 Instructions for Memory Construction
When processing long-context inputs, we split the content into chunks of a specified size and feed these chunks into the agent as memory. The agent can then extract relevant information from its memory based on the query to assist with query execution. This chunking approach helps organize and manage large amounts of contextual information, making retrieval and reasoning more efficient. In Figure 4, we provide several example instructions that require the agent to memorize the corresponding context.
B.2 Instructions for Long-Context Agents
In Figure 5, we provide the examples of instructions used on different of datasets. For some existing datasets, we adjust the prompt settings from previous work such as [15, 48]. For example, for the dataset Bench-QA and Bench-Sum, we also insert two answer examples as <demo> in the prompt to help the agent better understand the questions and standardize its outputs.
B.3 Instructions for RAG Agents
We provide examples of prompts used for the RAG based Agents in Figure 6. For this type of agent, after storing the input long context in memory, we use <question> as the memory retrieval query for most tasks. But for task RULER-NIAH-MQ, we use the entire question "What are all the special magic numbers for <question> mentioned in the memorized text?" as the query. And for Bench-Sum, we use the entire query without the <demo> for the memory retrieval.
| Agent Type | RULER-QA1 | RULER-QA2 | NIAH-MQ | Bench-QA | LME(S) | LME(S*) | EventQA |
| Long-Context Agents | |||||||
| GPT-4o | 72.0 | 51.0 | 25.0 | 55.4 | 61.4 | 32.0 | 77.2 |
| GPT-4o-mini | 64.0 | 43.0 | 22.8 | 44.9 | 55.6 | 30.7 | 59.0 |
| GPT-4.1-mini | 83.0 | 66.0 | 94.8 | 45.8 | 61.4 | 55.7 | 82.6 |
| Gemini-2.0-Flash | 87.0 | 59.0 | 83.8 | 53.2 | 52.6 | 47.0 | 67.2 |
| Claude-3.7-Sonnet | 77.0 | 53.0 | 38.0 | 50.6 | 59.0 | 34.0 | 74.6 |
| GPT-4o-mini | 64.0 | 43.0 | 22.8 | 44.9 | 55.6 | 30.7 | 59.0 |
| Simple RAG Agents | |||||||
| BM25 | 68.0 | 54.0 | 100.0 | 45.6 | 55.2 | 45.3 | 74.6 |
| Embedding RAG Agents | |||||||
| Contriever | 22.0 | 31.0 | 2.5 | 38.1 | 32.8 | 15.7 | 66.8 |
| Text-Embed-3-Small | 60.0 | 44.0 | 7.2 | 44.4 | 49.0 | 48.3 | 63.0 |
| Text-Embed-3-Large | 54.0 | 44.0 | 19.5 | 50.1 | 44.6 | 52.3 | 70.0 |
| NV-Embed-v2 | 90.0 | 67.0 | 73.5 | 51.4 | 45.4 | 55.0 | 72.8 |
| Structure-Augmented RAG Agents | |||||||
| RAPTOR | 29.0 | 38.0 | 15.8 | 31.3 | 38.8 | 34.3 | 45.8 |
| GraphRAG | 47.0 | 47.0 | 38.3 | 35.8 | 39.2 | 35.0 | 34.4 |
| HippoRAG-v2 | 76.0 | 66.0 | 67.5 | 45.7 | 44.2 | 50.7 | 67.6 |
| Mem0 | 24.0 | 32.0 | 4.8 | 22.4 | 45.0 | 36.0 | 37.5 |
| Cognee | 31.0 | 26.0 | 4.0 | 19.7 | 31.3 | 29.3 | 26.8 |
| Agentic Memory Agents | |||||||
| Self-RAG | 35.0 | 42.0 | 8.0 | 28.5 | 23.4 | 25.7 | 31.8 |
| MemGPT | 41.0 | 38.0 | 8.8 | 20.8 | 41.4 | 32.0 | 26.2 |
| Agent Type | BANKING | CLINC | NLU | TREC C | TREC F | Recom | Bench-Summ | FactCon-SH | FactCon-MH |
| Long-Context Agents | |||||||||
| GPT-4o | 96.0 | 96.0 | 90.0 | 87.0 | 69.0 | 12.3 | 32.2 | 60.0 | 5.0 |
| GPT-4o-mini | 93.0 | 93.0 | 87.0 | 73.0 | 66.0 | 15.1 | 28.9 | 45.0 | 5.0 |
| GPT-4.1-mini | 93.0 | 82.0 | 85.0 | 68.0 | 50.0 | 16.7 | 41.9 | 36.0 | 5.0 |
| Gemini-2.0-Flash | 91.0 | 90.0 | 84.0 | 88.0 | 67.0 | 8.7 | 23.9 | 30.0 | 3.0 |
| Claude-3.7-Sonnet | 97.0 | 98.0 | 86.0 | 87.0 | 79.0 | 18.3 | 52.5 | 43.0 | 2.0 |
| GPT-4o-mini | 93.0 | 93.0 | 87.0 | 73.0 | 66.0 | 15.1 | 28.9 | 45.0 | 5.0 |
| Simple RAG Agents | |||||||||
| BM25 | 89.0 | 89.0 | 84.0 | 62.0 | 53.0 | 13.6 | 20.9 | 56.0 | 3.0 |
| Embedding RAG Agents | |||||||||
| Contriever | 89.0 | 88.0 | 80.0 | 55.0 | 41.0 | 15.2 | 21.2 | 18.0 | 7.0 |
| Text-Embed-3-Small | 88.0 | 89.0 | 83.0 | 54.0 | 36.0 | 15.3 | 25.7 | 28.0 | 3.0 |
| Text-Embed-3-Large | 90.0 | 91.0 | 80.0 | 55.0 | 46.0 | 16.2 | 21.6 | 28.0 | 4.0 |
| NV-Embed-v2 | 88.0 | 89.0 | 82.0 | 40.0 | 48.0 | 13.5 | 20.7 | 55.0 | 6.0 |
| Structure-Augmented RAG Agents | |||||||||
| RAPTOR | 78.0 | 75.0 | 73.0 | 48.0 | 23.0 | 12.3 | 13.4 | 14.0 | 1.0 |
| GraphRAG | 64.0 | 54.0 | 49.0 | 24.0 | 6.0 | 9.8 | 0.4 | 14.0 | 2.0 |
| HippoRAG-v2 | 81.0 | 86.0 | 73.0 | 38.0 | 29.0 | 10.2 | 14.6 | 54.0 | 5.0 |
| Mem0 | 5.0 | 4.0 | 1.0 | 6.0 | 1.0 | 10.0 | 0.8 | 18.0 | 2.0 |
| Cognee | 34.0 | 42.0 | 42.0 | 41.0 | 18.0 | 10.1 | 2.3 | 28.0 | 3.0 |
| Agentic Memory Agents | |||||||||
| Self-RAG | 19.0 | 13.0 | 6.0 | 15.0 | 5.0 | 12.8 | 0.9 | 19.0 | 3.0 |
| MemGPT | 89.0 | 83.0 | 79.0 | 56.0 | 31.0 | 14.0 | 2.5 | 28.0 | 3.0 |
Appendix C Detailed Experimental Results
In this section, we provide detailed versions of the results presented in the main text.
C.1 Detailed Results on AR
In Table 7, we present the detailed results for each agent on every dataset. For AR tasks, using Simple RAG Agents equipped with retrievers like BM25 can significantly improve performance compared to the backbone model. This is because the GPT-4o-mini is limited by its 128K context length, which restricts the amount of information it can process at once. Meanwhile, the overall performance of Embedding RAG Agents surpasses that of both Structure-Augmented RAG Agents and Agentic Memory Agents. This advantage is primarily attributed to the use of dense retrieval in Embedding RAG Agents. It enables the extraction of longer contextual information from memory. As a result, Embedding RAG Agents are able to provide richer and more comprehensive context for tasks.
C.2 Detailed Results on TTL, LRU and CR
We give detailed results on each dataset in Table 7. For all three types of tasks, RAG-based agents generally underperform compared to their respective GPT-4o-mini backbones. This observation highlights certain limitations inherent to the RAG approach. For instance, in TTL tasks, RAG-based methods often struggle to accurately retrieve context from memory that is closely associated with the input. In LRU tasks, these methods face challenges in achieving a comprehensive understanding of long contexts. Furthermore, for CR tasks—especially the multi-hop variants—effective handling requires strong reasoning and information extraction capabilities, which remain beyond the reach of most current agents.
C.3 Detailed Results on Ablation Study
In this section, we introduce the detailed results on the ablation study on different chunk sizes, retrieve number, context length and computation latency.
In Table 8 and 11, we compare the RAG-based Agents on different chunk sizes and datasets. We selected chunk sizes from the two sets {512, 4096} and {512, 1024, 2048, 4096}. For some datasets composed of synthetic text, such as RULER-QA, using a smaller chunk size generally helps RAG Agents and Agentic Memory Agents achieve better test performance. However, for datasets composed of continuous text, such as Bench-QA, since the retrieval number remains unchanged, reducing the chunk size does not lead to performance improvement.
In Table 11, we evaluate the selected RAG-based Agents on five datasets. We choose different TopK ranging from {2, 5, 10}. We find that for the AR series of tasks, increasing the retrieve number (TopK) leads to a significant improvement in performance. However, for the TTL series of tasks, the performance gains from increasing TopK are much less pronounced.
In Table 11, we report the performances of different agents when scaling the input length. We measure the average context length via the tokenizer of GPT-4o-mini and here 1K is 1024. For Long-Context Agents, tasks in the AR series generally achieve satisfactory performance at relatively small context lengths (e.g., around 50K tokens). However, as the context length increases, the performance of these agents declines accordingly. In contrast, for the RAG-based agents Mem0 and Cognee, their performance is significantly lower than that of their backbone, GPT-4o-mini, even when the context length is relatively small.
In Table 12 and Table 14, we provide the computational latency on different agents. We choose two chunk sizes {512, 4096} and two datasets {RULER-QA2, LME(S*)}. The experimental results demonstrate that, for most agents, selecting a smaller chunk size leads to significantly higher computational latency. For example, in the case of Cognee, the computational latency at a chunk size of 512 is nearly 8 to 10 times greater than that at a chunk size of 4096.
| NIAH-MQ | Bench-QA | LME(S*) | Event-QA | FactCon-SH | FactCon-MH | |||||||
| 512 | 4096 | 512 | 4096 | 512 | 4096 | 512 | 4096 | 512 | 4096 | 512 | 4096 | |
| Simple RAG Agents | ||||||||||||
| BM25 | 100 | 95.5 | 32.9 | 45.6 | 45.3 | 48.3 | 69.4 | 74.6 | 56.0 | 44.0 | 3.0 | 5.0 |
| Embedding RAG Agents | ||||||||||||
| Contriever | 2.5 | 8.8 | 28.5 | 38.1 | 15.7 | 19.0 | 62.8 | 66.8 | 18.0 | 25.0 | 7.0 | 5.0 |
| Text-Embed-3-Small | 7.2 | 12.3 | 42.4 | 44.4 | 48.3 | 39.0 | 60.8 | 63.0 | 28.0 | 21.0 | 3.0 | 4.0 |
| Text-Embed-3-Large | 19.5 | 13.5 | 42.3 | 50.1 | 52.3 | 39.3 | 62.0 | 70.0 | 28.0 | 22.0 | 4.0 | 3.0 |
| NV-Embed-v2 | 73.5 | 31.8 | 40.7 | 51.4 | 55.0 | 43.0 | 72.7 | 72.8 | 55.0 | 42.0 | 6.0 | 6.0 |
| Structure-Augmented RAG Agents | ||||||||||||
| RAPTOR | 15.8 | 4.5 | 21.3 | 31.3 | 34.3 | 31.7 | 43.6 | 45.8 | 14.0 | 19.0 | 1.0 | 2.0 |
| GraphRAG | 38.3 | 8.0 | 35.2 | 35.8 | 35.0 | 36.7 | 33.8 | 34.4 | 14.0 | 10.0 | 2.0 | 3.0 |
| HippoRAG-v2 | 67.5 | 23.3 | 34.5 | 45.7 | 50.7 | 37.3 | 67.5 | 67.6 | 54.0 | 29.0 | 5.0 | 3.0 |
| Agentic Memory Agents | ||||||||||||
| Self-RAG | 8.0 | 7.0 | 27.7 | 28.5 | 25.7 | 23.0 | 30.2 | 31.8 | 19.0 | 14.0 | 3.0 | 2.0 |
| MemGPT | 8.8 | 3.5 | 23.5 | 20.8 | 32.0 | 28.0 | 26.2 | 25.8 | 28.0 | 13.0 | 3.0 | 3.0 |
| RULER-QA-1 | RULER-QA-2 | Bench-Sum | ||||||||||
| 512 | 1024 | 2048 | 4096 | 512 | 1024 | 2048 | 4096 | 512 | 1024 | 2048 | 4096 | |
| BM25 | 68.0 | 67.0 | 68.0 | 66.0 | 54.0 | 51.0 | 52.0 | 56.0 | 11.5 | 13.2 | 19.2 | 20.9 |
| NVEmbed-v2 | 90.0 | 80.0 | 57.0 | 57.0 | 67.0 | 59.0 | 52.0 | 39.0 | 11.6 | 13.0 | 16.8 | 20.7 |
| HippoRAG-v2 | 76.0 | 70.0 | 57.0 | 49.0 | 66.0 | 63.0 | 51.0 | 38.0 | 4.6 | 6.0 | 10.5 | 14.6 |
| MemGPT | 41.0 | 32.0 | 24.0 | 27.0 | 38.0 | 33.0 | 37.0 | 35.0 | 1.2 | 1.8 | 4.2 | 2.5 |
| RULER | NIAH | Bench-QA | EventQA | TTL (MCC) | |||||||||||
| R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | R=2 | R=5 | R=10 | |
| BM25 | 49.5 | 59.5 | 61.0 | 34.5 | 57.0 | 95.5 | 26.7 | 38.3 | 45.6 | 66.6 | 71.2 | 74.6 | 67.8 | 75.4 | 74.6 |
| Contriever | 26.0 | 38.0 | 41.0 | 0.5 | 5.3 | 8.8 | 23.8 | 37.1 | 38.1 | 54.4 | 66.8 | 56.0 | 63.0 | 70.0 | 70.6 |
| Text-Embed-3-Large | 32.5 | 33.5 | 36.5 | 5.0 | 9.3 | 13.5 | 34.8 | 41.9 | 50.1 | 51.8 | 62.4 | 70.0 | 59.4 | 69.4 | 72.4 |
| NV-Embed-v2 | 37.0 | 43.5 | 48.0 | 17.8 | 26.8 | 31.8 | 42.8 | 48.1 | 51.4 | 59.4 | 68.4 | 72.8 | 63.8 | 69.4 | 68.8 |
| RAPTOR | 21.0 | 19.5 | 23.5 | 4.3 | 4.5 | 4.3 | 30.9 | 30.4 | 31.3 | 45.8 | 41.8 | 40.4 | 56.3 | 59.4 | 57.4 |
| HippoRAG-v2 | 38.0 | 42.5 | 43.5 | 16.5 | 23.3 | 23.3 | 35.9 | 45.3 | 45.7 | 58.8 | 67.6 | 67.4 | 58.8 | 61.4 | 61.4 |
| Self-RAG | 29.5 | 32.5 | 38.5 | 4.0 | 6.3 | 7.0 | 21.0 | 23.9 | 28.5 | 28.2 | 30.6 | 31.8 | 9.0 | 11.6 | 11.6 |
| RULER | NIAH | EventQA | FactCon-SH | FactCon-MH | |||||||||||
| 51K | 104K | 304K | 55K | 117K | 448K | 51K | 108K | 534K | 32K | 64K | 262K | 32K | 64K | 262K | |
| GPT-4o | 81.5 | 76.0 | 61.5 | 100 | 100 | 25.0 | 96.8 | 94.0 | 77.2 | 88.0 | 85.0 | 60.0 | 10.0 | 13.0 | 5.0 |
| GPT-4o-mini | 71.0 | 68.5 | 53.5 | 99.5 | 99.0 | 22.8 | 90.2 | 85.8 | 59.0 | 63.0 | 58.0 | 45.0 | 10.0 | 5.0 | 5.0 |
| GPT-4.1-mini | 82.5 | 80.5 | 74.5 | 99.5 | 99.0 | 94.8 | 97.0 | 93.8 | 82.6 | 82.0 | 72.0 | 36.0 | 7.0 | 9.0 | 5.0 |
| Gemini-2.0-Flash | 80.5 | 74.0 | 73.0 | 87.8 | 93.3 | 83.8 | 93.4 | 88.6 | 67.2 | 49.0 | 62.0 | 30.0 | 7.0 | 9.0 | 3.0 |
| Claude-3.7-Sonnet | 78.5 | 70.5 | 65.0 | 99.5 | 100 | 38.0 | 96.6 | 95.2 | 74.6 | 46.0 | 45.0 | 43.0 | 2.0 | 2.0 | 0.0 |
| Mem0 | 28.5 | 27.0 | 28.0 | 5.5 | 5.0 | 4.8 | 60.8 | 47.0 | 40.2 | 22.0 | 8.0 | 18.0 | 3.0 | 2.0 | 2.0 |
| Cognee | 37.0 | 40.0 | 33.5 | 17.5 | 14.3 | 4.0 | 53.4 | 39.0 | 26.8 | 39.0 | 31.0 | 28.0 | 4.0 | 5.0 | 3.0 |
| RULER-QA2 | LME (S*) | |
| GPT-4o | 17.0 | 20.1 |
| GPT-4o-mini | 4.9 | 5.4 |
| GPT-4.1-mini | 9.0 | 7.4 |
| Gemini-2.0-Flash | 12.4 | 10.1 |
| Claude-3.7-Sonnet | 23.3 | 22.7 |
| RULER-QA2 | LME (S*) | |||||||
| 512 | 4096 | 512 | 4096 | |||||
| M.C. | Q.E. | M.C. | Q.E. | M.C. | Q.E. | M.C. | Q.E. | |
| BM25 | 0.12 | 0.47 | 0.11 | 1.7 | 0.09 | 1.1 | 0.08 | 1.9 |
| Contriever | 7.4 | 0.59 | 1.7 | 2.0 | 6.9 | 0.92 | 1.6 | 1.9 |
| Text-Embed-3-Large | 6.1 | 0.46 | 5.0 | 1.7 | 6.5 | 0.62 | 5.8 | 1.8 |
| NV-Embed-v2 | 102 | 0.63 | 47.0 | 1.8 | 85.1 | 1.0 | 38.8 | 1.7 |
| RAPTOR | 193 | 0.41 | 161 | 0.67 | 108 | 0.60 | 104 | 0.53 |
| GraphRAG | 97.8 | 12.8 | 91.9 | 10.9 | 149 | 7.0 | 88.8 | 7.8 |
| HippoRAG-v2 | 1089 | 0.71 | 380 | 1.71 | 544 | 1.5 | 188 | 3.5 |
| Mem0 | 10804 | 0.79 | 1334 | 0.65 | 18483 | 1.6 | 2946 | 1.7 |
| Cognee | 11890 | 58.7 | 1185 | 4.8 | 4728 | 7.7 | 738 | 4.1 |
| Self-RAG | 11.4 | 3.1 | 8.1 | 2.4 | 5.3 | 0.82 | 5.2 | 1.0 |
| MemGPT | 433 | 9.4 | 101 | 10.5 | 392 | 11.7 | 85.5 | 12.3 |
| Task | Max Output Tokens |
| RULER-QA | 50 |
| RULER-NIAH-MQ | 100 |
| Bench-QA | 10 |
| LongMemEval | 100 |
| EventQA | 40 |
| ICL_Five | 20 |
| Movie Recommendation | 300 |
| Bench-Sum | 1,200 |
| FactConsolidation | 10 |
Appendix D Experimental Settings
In this section, we present the experimental settings.
D.1 Max Output Tokens
We provide the token number limitation for each task in Table 14.
D.2 Settings of the RAG Agents
For the embedding model selection in Structure-Augmented RAG Agents and Agentic Memory Agents, most approaches utilize OpenAI’s embedding models, such as Text-Embed-3-Small. While for the HippoRAG-v2 method, we follow the same experimental setting as in Gutiérrez et al. [12], employing the NV-Embed-v2 model.
We implement three open-sourced memory agents in our main experiments. (1) For Mem0, we use memory.add() function to add the message with the content from each context chunk into the agent’s memory repository during memory consolidation. During query execution, the relevant memory elements are retrieved through memory.search() function. The retrieved memories are then integrated into the query before being processed by the GPT-4o-mini backbone model to complete the requested tasks. (2) For MemGPT, we employ the insert_passage() function during the memory consolidation phase to inject long context chunks into the Archival Memory structure. During query execution, this agent processes requests via the send_message() function which generates appropriate responses based on the archived information. (3) For Cognee, we utilize the cognee.add() and cognee.cognify() functions to construct the memory graph from input chunks wherein the memory consolidation phase. During query execution, the cognee.search() function is used to retrieve contextually relevant information from the memory graph based on the input query.
D.3 Settings of the Chunk Size
We use smaller chunk size (512) for synthetic context used in AR and CR. For some tasks based on continuous text, such as Bench and EventQA, we used a larger chunk size (4096). For tasks such as MCC, Recom and LME(S), considering the characteristics of these tasks and the computational cost, we also chose a larger chunk size (4096). For the two memory construction methods that are more time-consuming, Mem0 and Cognee, we uniformly used a chunk size of 4096 across all datasets.
| Chunk Size | 512 | 4096 |
| RULER-QA, NIAH-MQ | Bench-QA, Bench-Sum | |
| Dataset | FactCon-SH, FactCon-MH | MCC, Recom |
| LME(S*) | EventQA, LME(S) |